# The Adaptivity Barrier in Batched Nonparametric Bandits: Sharp Characterization of the Price of Unknown Margin

### v1

@@ Abstract
We study batched nonparametric contextual bandits under a margin condition when the margin parameter $\alpha$ is unknown. To capture the statistical price of this ignorance, we introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle knowing $\alpha$. We show that the optimal regret inflation grows polynomial with the horizon $T$, with exponent precisely given by the value of a convex optimization problem involving the dimension, smoothness, and batch budget. Moreover, the minimizers of this optimization problem directly prescribe the batch allocation and exploration strategy of a rate-optimal algorithm. Building on this principle, we develop RoBIN (RObust batched algorithm with adaptive BINning), which achieves the optimal regret inflation up to logarithmic factors. These results reveal a new adaptivity barrier: under batching, adaptation to an unknown margin parameter inevitably incurs a polynomial penalty, sharply characterized by a variational problem. Remarkably, this barrier vanishes when the number of batches exceeds $\log\log T$; with only a doubly logarithmic number of updates, one can recover the oracle regret rate up to polylogarithmic factors.
@@ 1 Introduction
A central question in sequential decision making is the cost of adaptation: how much performance is lost when key complexity parameters are unknown. Nonparametric contextual bandits [47, 48, 40, 31] provide a canonical setting to study this question. In the fully online regime, the problem is well understood. Under smoothness and margin assumptions, algorithms that attain minimax-optimal regret can even adapt to an unknown margin parameter at no extra cost. In particular, the foundational work of Rigollet and Zeevi [40] and the ABSE policy of Perchet and Rigollet [31] demonstrate that margin adaptation comes at no statistical cost in the online regime.
@@
In many settings, including clinical trials, education, and digital platforms, fully online interaction is infeasible because of logistical, ethical, or computational constraints. Instead, data collection proceeds in a limited number of batches: actions are fixed for a group of covariates, feedback is revealed only at the end of the batch, and subsequent policies must adapt accordingly. Such batch constraints arise naturally in domains ranging from clinical trials and education to digital platforms with delayed feedback. While minimax-optimal rates have been established for batched nonparametric contextual bandits when the margin parameter is known, these procedures require oracle knowledge to tune batch sizes and exploration schedules [22]. This raises a fundamental question: What is the statistical price of not knowing the margin parameter when learning under batch constraints?
@@
This paper provides a sharp answer. We introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle who knows the true margin parameter. We show that the optimal regret inflation grows polynomial with the horizon $T$, with an exponent precisely characterized by a convex variational optimization problem that depends on the dimension, smoothness, and batch budget. Strikingly, the minimizers of this program also prescribe the batch allocation and exploration schedule of a rate-optimal algorithm, yielding matching upper and lower bounds up to logarithmic factors.
@@
A key corollary is the identification of an adaptivity barrier unique to batching. In the online regime, the margin parameter admits free adaptation, but once updates are limited, adaptation becomes inherently costly. We prove that the barrier vanishes when the number of batches exceeds order $\log\log T$; that is, with only doubly logarithmic many updates, one can match the oracle regret rate up to polylogarithmic factors. Conversely, when the batch budget grows more slowly than $\log\log T$, the regret inflation is unavoidably polynomial. This threshold cleanly delineates the transition between the regimes where batching constrains adaptation and where it does not.
@@
Conceptually, our analysis unifies statistical limits and algorithm design through a single variational object. The variational characterization exposes the precise dependence of adaptivity cost on dimension, smoothness, and batch budget, while its minimizers yield an explicit constructive principle for designing robust batched policies under unknown complexity.
Organization of the paper.
Section 2 introduces the problem setup and the regret inflation criterion. Section 3 presents the main results, including the variational characterization of regret inflation and the master theorem. Section 4 describes the optimal algorithm guided by the variational principle. Section 5 contains the proof of the lower bound. Section 6 reviews the related work. We conclude with a discussion of extensions and future directions in Section 7.
Notation.
@@ 2.1 Model and assumptions
are drawn i.i.d. from a distribution $P$, where the context $X_{t}\in\mathcal{X}:=[0,1]^{d}$ has distribution $P_{X}$. The rewards take values in $[0,1]$ and satisfy
@@
The grid $\Gamma$ may be chosen adaptively, meaning that the statistician can use all information up to $t_{i-1}$ to determine $t_{i}$. The statistician’s policy $\pi_{t}$ at time $t$ is allowed to depend on $\mathcal{H}_{t}$. The goal of the statistician is to design an $M$-batch policy $(\Gamma,\pi)$ that can compete with an oracle that knows the environment, i.e., the law $P$ of $(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)})$.
@@
We impose the following standard conditions in the nonparametric bandits literature [40, 31, 8, 22]:
@@
$\underline{c}\,r^{d}\;\leq\;P_{X}\!\big(B(x,r)\big)\;\leq\;\bar{c}\,r^{d},\qquad\forall x\in\mathrm{supp}(P_{X}).\forall r\in(0,1],$
@@
Assumption 3 pertains to the margin condition in nonparametric classification [30, 44, 3], and has been adapted to the bandit setup by [16, 40, 31]. The margin parameter governs the fundamental complexity of the problem. When $\alpha=0$, the margin assumption becomes vacuous, and the reward functions of the two arms can be arbitrarily close to each other, making it challenging to identify the optimal one. When $\alpha$ increases, the reward functions of the two actions exhibit strong separation over a region of high probability mass, and discerning the optimal action is less difficult.
The following proposition adapted from [31] depicts the interplay between the smoothness parameter $\beta$ and the margin parameter $\alpha$.
@@ 2.2 Oracle regret with known margin
$R_{T}^{\star}(\alpha)=\inf_{(\Gamma,\pi)}\;\sup_{P\in\mathcal{P}_{\alpha}}\;R_{T}(\Gamma,\pi;P),$ (2)
where $\mathcal{P}_{\alpha}$ denotes the class of environments with margin $\alpha$. It is known that the rate depends on the smoothness $\beta$, dimension $d$, margin $\alpha$, and batch budget $M$. Define
@@
When $\alpha=\infty$, we have $\gamma(\infty)=\infty$, and $h_{M}(\infty)=0$.
The optimal rates with known margin have been established in [22], which are stated in the following proposition.
@@
$c_{1}M^{-4}\cdot T^{\,h_{M}(\alpha)}\leq R_{T}^{\star}(\alpha)\leq c_{2}M(\log T)\cdot T^{\,h_{M}(\alpha)},$ (4)
@@
Although [22] established the minimax rate for the case $\alpha\leq 1/\beta$, we extend their result to $\alpha\in[0,d/\beta]\cup\{\infty\}$ for $d\geq 1$ and also make the dependence on $M$ clear. See Appendix A for the proof.
@@ 2.3 The regret inflation criterion
This ratio compares the regret of the adaptive policy to that of the oracle who knows $\alpha$. Our goal is to characterize the rate of the optimal regret inflation, i.e., $\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi)$, and its dependence on the batch budget $M$.
3 Sharp characterization of regret inflation
@@
3.1 A variational problem
Minimizing regret inflation can be formulated as a two-player zero-sum game. The learner commits to an $M$-batch policy $(\Gamma,\pi)$ without knowledge of the margin parameter, while nature selects a distribution $P$ of contexts and rewards consistent with some margin parameter $\alpha$. The payoff is the ratio between the learner’s regret and that of the oracle who knows $\alpha$.
Although both the learner’s strategy space (policies) and nature’s strategy space (distributions) are infinite-dimensional, the complexity of this game can be captured by a finite-dimensional reduction. The learner’s choice reduces to specifying a batch allocation across the $M$ updates, and the nature’s choice reduces to selecting a margin parameter $\alpha$ from the admissible range $\mathcal{K}$. We parameterize the batch schedule $\Gamma=\{t_{0}=0<t_{1}<\cdots<t_{M}=T\}$ by exponents $u_{i}\in[0,1]$, so that $t_{i}\approx T^{u_{i}}$. We refer to $\bm{u}=(u_{1},\dots,u_{M-1})$ as an exponent grid. This reduction yields a finite-dimensional convex optimization problem.
Formally, let $\bm{u}\in\mathcal{U}_{M}=\{\bm{u}\in\mathbb{R}^{M-1}:0\leq u_{1}\leq\cdots\leq u_{M-1}\leq 1\}$ be the grid choice of the learner, and let $\alpha\in\mathcal{K}$ be the margin parameter selected by nature. Define the payoff function
@@ 3.2 Optimal regret inflation
For now, we focus on the dependence on $T$, i.e., assuming that $M$ is a fixed constant. We see from Theorem 1 that the exponent $\psi_{M}^{\star}$, determined by the variational problem, exactly quantifies the statistical price of not knowing the margin, up to logarithmic factors. It is thus instrumental to understand the behavior of $\psi_{M}^{\star}$. It turns out we always have $\psi_{M}^{\star}>0$, indicating that the regret inflation is inevitably polynomial in the horizon $T$; this is formally shown in the following proposition.
@@ 3.3 Numerical illustrations
3.4 How many batches suffice for no regret inflation?
Across all configurations in the above experiment, $\psi_{M}^{\star}$ decreases with the number $M$ of batches. This phenomenon is not obvious from the definition of the variational objective in (8): while more batches allow finer batch schedule and hence a lower regret, the benchmark $h_{M}(\alpha)$ also shrinks, so the improvement is nontrivial to analyze.
We can establish the monotonicity for small $M$.
Proposition 4.
For $M\leq 4$, we have $\psi_{M+1}^{\star}<\psi_{M}^{\star}$.
See Section B.2.3 for the proof.
While we are not able to prove the strict monotonicity of the optimal exponent $\psi_{M}^{\star}$ for general $M$, we can obtain a crude order-wise bound on $\psi_{M}^{\star}$ in terms of the number $M$ of batches.
Proposition 5.
See Appendix B for the proof.
Corollary 1.
There exists a constant $c_{1}$ such that if $M\geq c_{1}\log\log T$, then
where $C$ is some constant independent of $T$.
Conversely, by the lower bound in Proposition 5, we also know that $M\asymp\log\log T$ is necessary for small regret inflation. That is, when $M\leq c_{2}\log\log T$ for some small constant $c_{2}$, the regret inflation is super-polylogarithmic in $T$.
Thus $M\asymp\log\log T$ marks a sharp phase transition between regimes with costly and cost-free margin adaptation.

### v2

@@ Abstract
We study batched nonparametric contextual bandits under a margin condition when the margin parameter $\alpha$ is unknown. To capture the statistical cost of this ignorance, we introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle knowing $\alpha$. We show that the optimal regret inflation grows polynomially with the horizon $T$, with exponent $\psi_{M}^{\star}$ given by the value of a convex optimization problem that depends on the dimension, smoothness, and number of batches $M$. Moreover, the minimizer of this optimization problem directly prescribes the batch allocation and exploration strategy of a rate-optimal algorithm. Building on this principle, we develop RoBIN (RObust batched algorithm with adaptive BINning), which achieves the optimal regret inflation up to polylogarithmic factors. These results reveal a new adaptivity barrier: under batching, adaptation to an unknown margin parameter inevitably incurs a polynomial penalty, sharply characterized by a variational problem. Remarkably, this barrier vanishes once the number of batches exceeds order $\log\log T$; with only a doubly logarithmic number of updates, one can recover the oracle regret rate up to polylogarithmic factors.
@@ 1 Introduction
A central question in sequential decision making is the cost of adaptation: how much performance is lost when key complexity parameters are unknown. Nonparametric contextual bandits provide a canonical setting to study this question (Woodroofe, 1979; Yang and Zhu, 2002; Rigollet and Zeevi, 2010; Perchet and Rigollet, 2013). In the fully online regime, the problem is well understood. Under smoothness and margin assumptions, algorithms that attain minimax-optimal regret can even adapt to an unknown margin parameter at no extra cost. In particular, the foundational work of Rigollet and Zeevi (2010) and the ABSE policy of Perchet and Rigollet (2013) demonstrate that margin adaptation comes at no statistical cost in the online regime.
@@
In many settings, including clinical trials, education, and digital platforms, fully online interaction is infeasible because of logistical, ethical, or computational constraints. Instead, data collection proceeds in a limited number of batches: actions are fixed for a group of covariates, feedback is revealed only at the end of the batch, and subsequent policies must adapt accordingly. While minimax-optimal rates have been established for batched nonparametric contextual bandits when the margin parameter is known, these procedures require oracle knowledge to tune batch sizes and exploration schedules (Jiang and Ma, 2025). This raises a fundamental question: What is the statistical price of not knowing the margin parameter when learning under batch constraints?
@@
This paper provides a sharp answer. We introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle who knows the true margin parameter. We show that the optimal regret inflation grows polynomially as $T^{\psi_{M}^{\star}}$ with the horizon $T$, with an exponent $\psi_{M}^{\star}$ precisely characterized by a convex variational optimization problem that depends on the dimension, smoothness, and batch budget. Strikingly, the minimizer of this program also prescribes the batch allocation and exploration schedule of a rate-optimal algorithm, yielding matching upper and lower bounds up to polylogarithmic factors.
@@
A key corollary of this result is the identification of an adaptivity barrier unique to the batch constraint. In the online regime, the margin parameter admits free adaptation, but when updates are limited, adaptation becomes inherently costly. We prove that the barrier vanishes when the number of batches exceeds order $\log\log T$; that is, with a doubly logarithmic number of updates, one can match the oracle regret rate up to polylogarithmic factors. Conversely, when the batch budget grows more slowly than $\log\log T$, the regret inflation is unavoidably polynomial. This threshold cleanly delineates the transition between the regimes where batching constrains adaptation and where it does not.
@@
Conceptually, our analysis unifies statistical limits and algorithm design through a single variational object. The variational characterization exposes the precise dependence of adaptivity cost on dimension, smoothness, and batch budget, while its minimizer yields an explicit constructive principle for designing robust batched policies under unknown complexity.
1.1 Related work
Contextual bandits.
The concept of contextual bandits was introduced by Woodroofe (1979). For linear contextual bandits, Auer (2002); Abbasi-Yadkori et al. (2011); Goldenshluger and Zeevi (2013); Bastani and Bayati (2020); Qian et al. (2023) established regret guarantees in both low- and high-dimensional settings. Meanwhile, modeling the mean reward function as a smooth function of the contexts was studied in Yang and Zhu (2002). Rigollet and Zeevi (2010) proved a minimax regret lower bound for this setup and designed an upper-confidence-bound-type (UCB-type) policy to attain the near-optimal rate. Perchet and Rigollet (2013) refined this result by proposing the Adaptively Binned Successive Elimination (ABSE) policy that can also adapt to the unknown margin parameter in the fully online setting. Additional insights in nonparametric contextual bandits were obtained in Qian and Yang (2016a); Reeve et al. (2018); Guan and Jiang (2018); Hu et al. (2022); Suk and Kpotufe (2021); Gur et al. (2022); Cai et al. (2024).
Margin condition in classification.
The margin condition originates from nonparametric classification, where it was studied by Audibert and Tsybakov (2007). This condition, often referred to as the Tsybakov margin condition, quantifies how well-separated the optimal decision boundary is and directly governs learning rates. Its adaptation to contextual bandits was initiated by Goldenshluger and Zeevi (2009); Rigollet and Zeevi (2010); Perchet and Rigollet (2013), who showed that the margin parameter $\alpha$ fundamentally shapes the complexity of bandit learning. In the online setting, adaptation to unknown $\alpha$ is possible without additional regret cost (Perchet and Rigollet, 2013). However, under the batch constraint, as explored in our work, such adaptivity becomes costly, giving rise to a new barrier.
Batch learning.
The multi-armed bandit problem under the batched setting was studied by Perchet et al. (2016); Gao et al. (2019). Batch learning in linear contextual bandits was studied by Han et al. (2020); Ren et al. (2020); Ruan et al. (2021) and Ren and Zhou (2023); Wang and Cheng (2020); Fan et al. (2023) further considered the problem with high-dimensional covariates. Jiang and Ma (2025); Arya (2025) studied the nonparametric contextual bandit problem under the batch constraint. Karbasi et al. (2021); Kalkanli and Ozgur (2021) developed batched Thompson sampling algorithms. Feng et al. (2022) considered the Lipschitz continuum-armed bandit problem under the batched setting. Further insights in batched bandits were developed in Zhang et al. (2020a); Jin et al. (2021a, b); Liu et al. (2025). Another related topic is online learning with switching costs (Cesa-Bianchi et al., 2013). Best arm identification with limited rounds of interaction has been studied by Tao et al. (2019). Reinforcement learning with low switching costs has been considered by Bai et al. (2019); Zhang et al. (2020b); Gao et al. (2021); Wang et al. (2021); Qiao et al. (2022).
Adaptation.
In the fully online setting, adaptivity to the margin parameter is feasible at no extra cost (Perchet and Rigollet, 2013). One might ask whether the same is true for the smoothness parameter. The answer is negative: even without the batch constraint, adaptation to smoothness is impossible (Locatelli and Carpentier, 2018; Gur et al., 2022; Cai and Pu, 2022). Minimax regret rates depend explicitly on the Hölder smoothness $\beta$, and no single procedure can achieve the optimal rate simultaneously across different values of $\beta$. This impossibility parallels classical results in nonparametric estimation and classification, where smoothness adaptation requires additional structure or necessarily incurs a penalty (Qian and Yang, 2016b; Gur et al., 2022; Cai et al., 2024). Thus, the margin parameter is the quantity of genuine interest: it admits free adaptation online, yet, becomes costly under the batch constraint.
1.2 Notation and paper organization
@@
In terms of paper organization, in Section 2, we introduce the problem setup and the regret inflation criterion. Section 3 presents the main results, including the variational characterization of regret inflation and the master theorem. Section 4 describes the optimal algorithm guided by the variational principle. Section 5 contains the proof of the lower bound. We conclude with a discussion of extensions and future directions in Section 6.
@@ 2.1 Model and assumptions
are drawn i.i.d. from a distribution $P$, where the context $X_{t}\in\mathcal{X}\coloneqq[0,1]^{d}$ follows a distribution $P_{X}$. The rewards take values in $[0,1]$ and satisfy
@@
The grid $\Gamma$ may be chosen adaptively, meaning that the statistician can use all information up to $t_{i-1}$ to determine $t_{i}$. The statistician’s policy $\pi_{t}$ at time $t$ may depend on $\mathcal{H}_{t}$. The goal is to design an $M$-batch policy $(\Gamma,\pi)$ that can compete with an oracle that knows the environment, i.e., the law $P$ of $(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)})$.
@@
We impose the following standard conditions in the nonparametric bandits literature (Rigollet and Zeevi, 2010; Perchet and Rigollet, 2013; Cai et al., 2024; Jiang and Ma, 2025):
@@
$\underline{c}\,r^{d}\;\leq\;P_{X}\!\big(B(x,r)\big)\;\leq\;\bar{c}\,r^{d},\qquad\forall x\in\mathrm{supp}(P_{X}),\forall r\in(0,1],$
@@
Assumption 3 pertains to the margin condition in nonparametric classification (Mammen and Tsybakov, 1999; Tsybakov, 2004; Audibert and Tsybakov, 2007), and has been adapted to the bandit setup by Goldenshluger and Zeevi (2009); Rigollet and Zeevi (2010); Perchet and Rigollet (2013). The margin parameter governs the fundamental complexity of the problem. When $\alpha=0$, the margin assumption becomes vacuous, and the reward functions of the two arms can be arbitrarily close to each other, making it challenging to identify the optimal one. When $\alpha$ increases, the reward functions of the two actions exhibit strong separation over a region of high probability mass, and discerning the optimal action is less difficult.
The following proposition adapted from Perchet and Rigollet (2013) depicts the interplay between the smoothness parameter $\beta$ and the margin parameter $\alpha$.
@@ 2.2 Oracle regret with known margin
$R_{T}^{\star}(\alpha)\coloneqq\inf_{(\Gamma,\pi)}\;\sup_{P\in\mathcal{P}_{\alpha}}\;R_{T}(\Gamma,\pi;P),$ (2)
where $\mathcal{P}_{\alpha}$ denotes the class of environments with margin $\alpha$. It is known that the rate depends on the dimension $d$, smoothness $\beta$, margin $\alpha$, and batch budget $M$. Define
@@
When $\alpha=\infty$, we have $\gamma(\infty)=\infty$, and $h_{M}(\infty)=0$. The optimal rates with known margin have been established in Jiang and Ma (2025), which are stated in the following proposition.
@@
$c_{1}M^{-4}T^{\,h_{M}(\alpha)}\leq R_{T}^{\star}(\alpha)\leq c_{2}M(\log T)T^{\,h_{M}(\alpha)},$ (4)
@@
Although Jiang and Ma (2025) established the minimax rate for the case $\alpha\leq 1/\beta$, we extend their result to $\alpha\in[0,d/\beta]\cup\{\infty\}$ for $d\geq 1$. Importantly, we make the dependence on $M$ explicit in the lower bound. See Appendix A for the proof.
@@ 2.3 The regret inflation criterion
This ratio compares the regret of the adaptive policy to that of the oracle who knows $\alpha$. Our notion of regret inflation parallels the risk inflation criterion in model selection for linear models (Foster and George, 1994). In that context, risk inflation measures the excess risk of a data-driven model selection rule relative to an oracle that knows the true model. Analogously, regret inflation quantifies the excess regret of an adaptive bandit policy relative to an oracle that knows the margin parameter. Both notions capture the statistical price of adaptation to unknown complexity.
Our goal is to characterize the rate of the optimal regret inflation,
$\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi),$
and its dependence on the batch budget $M$.
3 Sharp characterization of the optimal regret inflation
@@
3.1 Key variational problem
Minimizing regret inflation can be formulated as a two-player zero-sum game. The learner commits to an $M$-batch policy $(\Gamma,\pi)$ without knowledge of the margin parameter, while nature selects a distribution $P$ over contexts and rewards consistent with some margin parameter $\alpha$. The payoff is the ratio between the learner’s regret and that of the oracle who knows $\alpha$.
Although both the learner’s strategy space (policies) and nature’s strategy space (distributions) are infinite-dimensional, the complexity of this game can be captured by a finite-dimensional reduction. The learner’s choice reduces to specifying a batch allocation across the $M$ updates, and nature’s choice reduces to selecting a margin parameter $\alpha$ from the admissible range $\mathcal{K}$. We parameterize the batch schedule $\Gamma=\{t_{0}=0<t_{1}<\cdots<t_{M}=T\}$ by exponents $u_{i}\in[0,1]$, so that $t_{i}\approx T^{u_{i}}$. This exponent parameterization captures exponentially growing batch sizes, which balance exploration and exploitation under geometric time scales. We refer to $\bm{u}=(u_{1},\dots,u_{M-1})$ as an exponent grid. This reduction yields a finite-dimensional convex optimization problem.
Formally, let $\bm{u}\in\mathcal{U}_{M}\coloneqq\{\bm{u}\in\mathbb{R}^{M-1}:0\leq u_{1}\leq\cdots\leq u_{M-1}\leq 1\}$ be the grid choice of the learner, and let $\alpha\in\mathcal{K}$ be the margin parameter selected by nature. Define the payoff function
@@ 3.2 Optimal regret inflation
To better interpret Theorem 1, we treat $M$ as fixed and focus on how the regret inflation scales with the horizon $T$;the effect of varying $M$ will be analyzed in Section 3.4. Theorem 1 shows that the variational value $\psi_{M}^{\star}$ fully determines this dependence: up to logarithmic factors, any adaptive policy incurs inflation of order $T^{\psi_{M}^{\star}}$. If $\psi_{M}^{\star}$ is strictly positive, then Theorem 1 implies an unavoidable polynomial adaptation cost in $T$. The next subsection analyzes the structure of the variational problem and establishes that $\psi_{M}^{\star}$ is indeed strictly positive.
3.3 Structural properties of the variational problem
@@
Part 1 of Proposition 3 ensures the exponent is strictly positive, implying polynomial inflation (up to polylogs) in the fixed $M$ case via Theorem 1. In Part 2 of Proposition 3, the sequence $\{\alpha_{i}\}$ identifies the branches of $\Psi_{M}$ that are active at $\bm{u}^{\star}$. This structure is used in the lower-bound construction (Section 5) to select difficulty levels of the hard instances. See Appendix B for the proof.
3.4 How many batches suffice for no regret inflation?
When we allow the number $M$ of batches to increase with the horizon $T$, a natural question arises: how many batches are sufficient for no regret inflation?
The following proposition provides a crude order-wise bound on $\psi_{M}^{\star}$ in terms of the number $M$ of batches.
Proposition 4.
See Appendix B for the proof.
Corollary 1.
There exists a constant $c_{1}$ such that if $M\geq c_{1}\log\log T$, then
where $C$ is some constant independent of $T$.
Conversely, by the lower bound in Proposition 4, we also know that $M\asymp\log\log T$ is necessary for small regret inflation. That is, when $M\leq c_{2}\log\log T$ for some small constant $c_{2}$, the regret inflation is super-polylogarithmic in $T$. Together, $M\asymp\log\log T$ marks a sharp phase transition between regimes with costly and cost-free margin adaptation.